### v2

# Dynamical localization and eigenvalue asymptotics: long-range hopping lattice operators with electric field

Abstract
We prove that for polynomial long-range hopping lattice operators with uniform electric field under any bounded perturbation, the semi-uniform polynomial decay of the eigenfunctions is determined by the asymptotic behavior of the eigenvalues, and conversely. Consequently, we recover and refine recent results on power-law dynamical localization for this model. In this paper, to prove these results, we develop new arguments to deal with singularities in obtaining the semi-uniform polynomial decay of eigenfunctions, prove an independent result on the asymptotic behavior of eigenvalues for discrete operators, revisit a notion of Power-Law SULE, and utilize a known perturbation result on localization, proved through the KAM method. Unlike existing results in the literature, our approach does not rely on the specific form of the eigenfunctions, but rather the asymptotic behavior of the eigenvalues and the potential. It is worth underlining that our general results can be applied to other models such as Maryland-type potentials.
Keywords: Dynamical Localization, Discrete Spectrum, Spectral Theory, Quantum Dynamics.
AMS classification codes: 28A80 (primary), 42A85 (secondary).
1 Introduction
1.1 Contextualization
There is a vast literature concerning the dynamics of wave packet solutions $e^{-itH}\psi$ of the Schrödinger equation
$\begin{cases}\partial_{t}\psi=-iH\psi,~{}t\in\mathbb{R},\\ \psi(0)=\psi,~{}\psi\in\text{dom}H,\end{cases}$ (1)
where $H$ is a self-adjoint operator in $\ell^{2}(\mathbb{Z})$. Actually, the study of the dynamics $e^{-itH}\psi$ through the spectral properties of $H$ is a classical subject in Quantum Dynamics; for instance, see [6, 11, 13, 17, 23, 25, 27] and references therein. In this context, a dynamical quantity to probe the asymptotic behavior of $e^{-itH}\psi$ is the $q$-moment, $q>0$, of the position operator at time $t>0$, defined as
$\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-itH}\psi,\delta_{n}\rangle|^{2},$
where $\delta_{n}=(\delta_{nj})_{j\in\mathbb{Z}}$ is the canonical basis of $\ell^{2}(\mathbb{Z})$. These quantities characterize the spreading of the wave packet $e^{-itH}\psi$.
In recent years, there has been substantial research activity particularly dedicated to the study of the phenomenon of dynamical localization [2, 5, 7, 8, 10, 15, 19, 20, 21, 22, 26], measured through the moments of order $q>0$. $H$ exhibits dynamical localization if, for all initial conditions $\delta_{k}$, each moment is uniformly bounded in time, that is, for each $q>0$ and each $\delta_{k}$,
$\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-itH}\delta_{k},\delta_{n}\rangle|^{2}<+\infty.$
This is the standard strong notion of absence of transport for discrete operators. Due to the RAGE Theorem [6], dynamical localization requires pure point spectrum; however, it is possible to have systems with pure point spectrum without dynamical localization [11, 13]. One way to establish such a property is through the control of eigenfunctions, referred to as Semi-Uniform Localization of Eigenfunctions (SULE) [11, 27].
Let $H$ be a self-adjoint operator acting in $\ell^{2}(\mathbb{Z})$. We say that $H$ exhibits SULE if there exists an orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $H$, $\{\phi_{m}\}_{m\in\mathbb{Z}}$, that is, $H\phi_{m}=\lambda_{m}\phi_{m},\,m\in\mathbb{Z}$, for which there exist $\alpha>0$ and, for each eigenfunction $\phi_{m}$, an $j_{m}\in\mathbb{Z}$ such that, for any $\delta>0$,
$|(\phi_{m})(n)|\leq\gamma_{\delta}e^{\delta|j_{m}|-\alpha|n-j_{m}|},$
where $\gamma_{\delta}>0$ is uniform in $m$ and $n$. Consequently, if $H$ satisfies SULE, then $H$ has dynamical localization (Theorem 7.5 in [11]).
An important model in quantum dynamics that satisfies SULE is the discrete one-dimensional Schrödinger operator with uniform electric field of strength $E=1$ [10, 18]:
$\text{dom }H_{0}:=\{u\in\ell^{2}(\mathbb{Z}):\sum_{n\in\mathbb{Z}}|n|^{2}|u(n)|^{2}<+\infty\},$
$(H_{0}u)(n)=u(n-1)+u(n+1)+nu(n),\quad n\in\mathbb{Z}.$
de Oliveira and Pigossi [10] proved that $H_{0}$ exhibits semi-uniform decay of eigenfunctions. For this model, eigenvalues and eigenfunctions were explicitly computed using Fourier series in [10], revealing that the spectrum consists of the set of integers. In particular, it was established that $H_{0}$ exhibits SULE (Semi-Uniform Localization of Eigenfunctions) and, consequently, dynamical localization. Moreover, through an iterative matrix diagonalization process originally developed by Kolmogorov, Arnold, and Moser (KAM), it was shown that for sufficiently small perturbations, the semi-uniform decay of eigenfunctions persists, thereby establishing the persistence of dynamical localization. In this case, it was also shown that the asymptotic behavior of the eigenvalues is preserved. In fact, the following result was established.
Theorem 1.1 (Theorem 4.2 in [10]).
To every $b\in\ell^{\infty}(\mathbb{Z})$, the operator $H_{0}+b$ has purely simple discrete spectrum; if $||b||_{\infty}$ is small enough, then there exists an orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $H_{0}+b,$$\{\phi_{m}\}_{m\in\mathbb{Z}},$ that satisfies
$|\phi_{m}(n)|\leq\gamma\,e^{-|m-n|},n,m\in\mathbb{Z},$
for some $\gamma>0$; in particular, $H_{0}+b$ has dynamical localization. Moreover, there exists a $\gamma^{\prime}>0$ such that its eigenvalues satisfy
$|\lambda_{n}-n|\leq\gamma^{\prime},n\in\mathbb{Z}.$
Remark 1.1.
Actually, de Oliveira and Pigossi in [8], using an iterative matrix diagonalization process (KAM), proved the persistence of the semi-uniform decay of eigenfunctions for sufficiently small perturbations of self-adjoint operators with discrete spectrum that possess this property, since the modulus of the difference between the eigenvalues of the original operator is lower bounded by a uniform constant. For example, in the case of $H_{0}$, one has $|\lambda_{m}-\lambda_{n}|=|m-n|\geq 1$ for $m\neq n$.
Still with respect to models with uniform electric field, consider the polynomial long-range hopping lattice operators.
Let $(a_{n})_{n\in\mathbb{Z}}$ in ${\mathbb{C}}^{\mathbb{Z}}$ such that $a(0)=0$, $a(m)=a^{*}(-m)$, $m\in\mathbb{Z}$, and $a\in\ell^{1}_{r}(\mathbb{Z})$, $r\geq 0$, that is,
$||a||_{r}=\sum_{m\in\mathbb{Z}}|a(m)||m|^{r}<+\infty.$
Let $T_{0}$ given by
$\text{dom }T_{0}:=\{u\in\ell^{2}(\mathbb{Z}):\sum_{n\in\mathbb{Z}}|n|^{2}|u(n)|^{2}<+\infty\},$
$(T_{0}u)(n)=(T_{a}u)(n)+nu(n)=\sum_{m\in\mathbb{Z}}a(n-m)u(m)+nu(n),\quad n\in\mathbb{Z}.$
Recently, Sun and Wang in [26] established the following results regarding localization for this model with uniform electric field.
Theorem 1.2 (Theorem 1.1 in [26]).
Fix $r>1$ and let $a\in\ell^{1}_{r}(\mathbb{Z})$. To every $b\in\ell^{\infty}(\mathbb{Z})$, $T_{0}+b$ has purely simple discrete spectrum. If
$\frac{1}{2}<s<r-\frac{1}{2}$
and $||b||_{\infty}$ is small enough, then there exists an orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $T_{0}+b,$$\{\phi_{m}\}_{m\in\mathbb{Z}},$ that satisfies
$|\phi_{m}(n)|\leq\frac{\gamma_{s,r}}{|n-m|^{s}},$
where $\gamma_{s,r}$ is a constant that depends only on $s$ and $||a||_{r}$.
Theorem 1.3 (Theorem 1.2 in [26]).
Fix $r>1$. Let $T_{0}+b$ before. If
$10\leq q<s-\frac{1}{2}<r-$ 1 (2)
and $||b||_{\infty}$ is small enough, then for any $\psi\in\ell^{2}_{q}(\mathbb{Z})$, one has
$\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{2q}|\langle e^{-i(T_{0}+V)t}\psi,\delta_{n}\rangle|^{2}<+\infty.$
Moreover, there exists a $\gamma^{\prime}>0$ such that the eigenvalues satisfy
$|\lambda_{n}-n|\leq\gamma^{\prime},n\in\mathbb{Z}.$
Remark 1.2.
  1. i)
    If $q>\frac{1}{2}$, the assumption (2) in Theorem 1.2 can be replaced by $\frac{1}{2}<q\leq s<r-\frac{1}{2}$.
  2. ii)
    $H_{0}$ is a particular case of a long-range hopping lattice operator. Thus, the results of Sun and Wang presented above extend the result of Pigossi and de Oliveira. However, it is worth mentioning that, although Sun and Wang obtained uniform power-law dynamical localization, this differs from the result of Pigossi and Oliveira, where exponential localization directly leads to dynamical localization without restrictions on $q>0$. We note that Pigossi and de Oliveira in [7] have also recently proven exponential dynamical localization for polynomial long-range hopping lattice operators when $a(m)$ has finite support.
Unlike ergodic Schrödinger operators (such as the Anderson model), where small divisors or resonances may occur (see [2, 5, 12, 19, 20] and references therein), the model considered here lacks resonances, making it possible to establish rigidity results that relate the spectral structure to the decay of eigenfunctions.
Here we prove that for polynomial long-range hopping lattice operators with uniform electric field under any bounded perturbation, the semi-uniform polynomial decay of the eigenfunctions is determined by the asymptotic behavior of the eigenvalues, and conversely.
We shall prove the following
Theorem 1.4.
Let $T_{0}$ be as before and let $T_{0}+b$, $b\in\ell^{\infty}(\mathbb{Z})$; so $T_{0}+b$ has purely discrete spectrum. Are equivalent
  1. i)
    there exists a $\gamma^{\prime}>0$ such that the eigenvalues satisfy
    $|\lambda_{n}-n|\leq\gamma^{\prime},n\in\mathbb{Z},$
  2. ii)
    if $a\in\ell^{1}_{r}(\mathbb{Z})$, $r\in\mathbb{N}\cup\{0\}$, then any orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $T_{0}+b,$$\{\phi_{m}\}_{m\in\mathbb{Z}},$ satisfies
    $|\phi_{m}(n)|\leq\frac{\gamma_{r}}{|n-m|^{r+1}},$
    where $\gamma_{r}$ is a constant that depends only on r and $||a||_{r}$,
Remark 1.3.
  1. i)
    Theorem 1.4 establishes that there exists a symbiosis between the semi-uniform polynomial decay of the eigenfunctions and the asymptotic behavior of the eigenvalues, which should be the same as that of the potential.
  2. ii)
    Note that, in item $i)$, the result holds for any orthonormal basis. Namely, here, we do not use a specific form for the eigenfunctions.
  3. iii)
    The decay condition
    $|\phi_{m}(n)|\leq\frac{\gamma}{|n-m|},\,\gamma>0,$
    is sufficient to ensure that $i)$ holds (see Theorem 2.2 ahead). We state the result in this form as it is more natural due to potential applications (see also Theorem 1.5 below).
We note that it is straightforward to use Fourier series to verify that the spectrum of $T_{0}$ is given by the integers. In this sense, the conclusions of item $ii)$ can be obtained directly for $b=0$ and later extended to sufficiently small perturbations using the KAM method. Consequently, the conclusions of item $i)$ also apply in this case. Namely, by combining Theorem 1.4 with a notion of Power-Law SULE (Lemma 3.1), we recover and refine the results discussed above on power-law dynamical localization for this model.
We shall prove the following
Theorem 1.5.
Let $T_{0}+b$ be as before. If $||b||_{\infty}$ is small enough, so,
  1. i)
    there exists a $\gamma^{\prime}>0$ such that the eigenvalues satisfy
    $|\lambda_{n}-n|\leq\gamma^{\prime},n\in\mathbb{Z},$
  2. ii)
    if $a\in\ell^{1}_{r}(\mathbb{Z})$, $r\in\mathbb{N}\cup\{0\}$, then any orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $T_{0}+b,$$\{\phi_{m}\}_{m\in\mathbb{Z}},$ satisfies
    $|\phi_{m}(n)|\leq\frac{\gamma_{r}}{|n-m|^{r+1}},$
    where $\gamma_{r}$ is a constant that depends only on r and $||a||_{r}$,
  3. iii)
    if $10<q<2r-$ 1, then for any $\psi=\delta_{k}$, one has
    $\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-i(T_{0}+b)t}\psi,\delta_{n}\rangle|^{2}<+\infty.$
Remark 1.4.
  1. i)
    Our main contribution in this paper is to show that there is a symbiotic relation between the structure of the spectrum/potential and the semi-uniform polynomial decay of eigenfunctions (Theorem 1.4) for models with electric field, which directly leads (Lemma 3.1) to power-law dynamical localization (Theorem 1.5). To the best of the present author’s knowledge, none of this has yet been detailed in the literature. It is worth underlining that our strategy for proving Theorem 1.5 is different from the proofs of Theorems 1.2 and 1.3 in [26]. In fact, here we develop a theory on the asymptotic behavior of eigenvalues of operators with potentials that grow according to a power-law (Theorems 2.1 and 2.2), combined with a notion of Power-Law SULE (Lemma 3.1), whereas in [26] the authors deal with Sobolev-type norms. It is worth mentioning that our notion of Power-Law SULE (Lemma 3.1) applies to other models, such as Maryland-type potentials [24] (see Lemma 3.1 ahead and Corollary 2.10 in [20]).
  2. ii)
    1) In [19], Shi introduces a new approach in the spirit of multi-scale analysis (MSA) of Frohlich-Spencer [12], developing an MSA approach that addresses power-law localization, eliminating the absolute continuity restrictions on the random distribution in Aizenman-Molchanov’s work [1]. 2) A subtle issue in the notion of SULE is that the very definition and the fact that $\{\phi_{m}\}_{m\in\mathbb{Z}}$ is an orthonormal basis ensure that the sequence $\{j_{m}\}_{m\in\mathbb{Z}}$ is not arbitrary, in the sense that $A_{0}=\sum_{m}e^{-\delta|j_{m}|}<+\infty$ (see (7.4) in [11]), which allows one to conclude that SULE implies dynamical localization. Namely, this is related to the fact that, in the case of SULE, the center of the eigenfunctions is $1$ (Theorem 7.1 in [11]). In this paper, we revisit a notion of Power-Law SULE and utilize asymptotic control of eigenvalues to obtain results on power-law dynamical localization inspired by these ideas. Namely, the main ingredients of the proofs of Theorems 1.4 and 1.5 are a notion of (Strong) Power-Law SULE (Lemma 3.1), an independent result on the asymptotic behavior of eigenvalues for unbounded discrete operators (Theorem 2.2) which is obtained from concentration results for orthonormal bases in $\ell^{2}(\mathbb{Z})$ and a known perturbation result (Theorem 1.1 in [8]) on dynamical localization proven through the KAM method.
  3. iii)
    Although the discrete spectrum of $T_{0}+b$ for any bounded potential $b$ provides strong evidence of dynamical localization, it remains an open problem whether dynamical localization holds for large $\|b\|_{\infty}$ or not.
  4. iv)
    It would be interesting to analyze the technique presented in this work in the context of time-periodic perturbations of models with uniform electric field, which were discussed in [10, 14].
1.2 Organization of text
In Section 2, we discuss the stability of the asymptotic behavior of eigenvalues for perturbations of unbounded discrete self-adjoint operators, under the hypothesis of semi-uniform polynomial decay of eigenfunctions (Theorems 2.1 and 2.2), and we prove Theorem 1.4. Section 3 addresses how a variation of SULE implies a variation of dynamical localization (Lemma 3.1), and then we prove Theorem 1.5. In Section 4, we prove concentration results for orthonormal bases in $\ell^{2}(\mathbb{Z})$, assuming a hypothesis of semi-uniform polynomial decay, which will be used in the proofs of Theorems 2.1 and 2.2. Section 5 presents the proofs of Theorems 2.1 and 2.2. In Appendix A, we revisit a property of the discrete spectrum.
Some words about notation: $\ell^{2}(\mathbb{Z})$ represents the space of square-summable sequences indexed by the integers $\mathbb{Z}$, that is,
$\ell^{2}(\mathbb{Z})=\left\{\psi:\mathbb{Z}\longrightarrow\mathbb{C}\;:\;\sum_{n\in\mathbb{Z}}|\psi(n)|^{2}<+\infty\right\}.$
This space is equipped with the usual inner product given by
$\langle\psi,\phi\rangle=\sum_{n\in\mathbb{Z}}\overline{\psi(n)}\phi(n),\quad\psi,\phi\in\ell^{2}(\mathbb{Z}).$
The induced norm is defined as $\|\psi\|_{2}=\sqrt{\langle\psi,\psi\rangle}$. $\{\delta_{n}:=(\delta_{nj})_{j\in\mathbb{Z}}\}$ will always denote the canonical basis in $\ell^{2}(\mathbb{Z})$. Let $H$ be a self-adjoint operator defined in $\ell^{2}(\mathbb{Z})$. The resolvent set of $H$, denoted by $\rho(H)$, is the set of values $\lambda\in\mathbb{C}$ for which $(H-\lambda I)^{-1}$ exists and is a bounded operator. The spectrum of $H$ is defined as
$\sigma(H)=\mathbb{C}\setminus\rho(H).$
In this paper, $A$ will always denote a bounded self-adjoint operator on $\ell^{2}(\mathbb{Z})$ and $I$ will denote the identity operator. $\|f\|_{\infty}$ denotes the norm of a sequence $f=(f(n))_{n\in\mathbb{Z}}$ in $\ell^{\infty}(\mathbb{Z})$, which is given by
$\|f\|_{\infty}=\sup_{n\in\mathbb{Z}}|f(n)|.$
Finally, $\ell^{2}(\mathbb{Z}_{+})$ and $\ell^{2}(\mathbb{Z}_{-})$ are defined as follows:
$\ell^{2}(\mathbb{Z}_{\mp})=\left\{\psi:\mathbb{Z}_{\mp}\longrightarrow\mathbb{C}\;:\;\sum_{n\in\mathbb{Z}_{\mp}}|\psi(n)|^{2}<+\infty\right\}.$
2 Eigenvalue asymptotics and decay of eigenfunctions
2.1 Eigenvalue asymptotics
Let $A$ be a bounded operator in $\ell^{2}(\mathbb{Z})$ and let $V:\mathbb{Z}\longrightarrow\mathbb{R}$. Consider
$\text{dom }H:=\{u\in\ell^{2}(\mathbb{Z}):\sum_{n\in\mathbb{Z}}|V(n)|^{2}|u(n)|^{2}<+\infty\},$
$(Hu)(n)=(Au)(n)+V(n)u(n),\quad n\in\mathbb{Z};$
so, by the Kato-Rellich Theorem [6], $H$ is self-adjoint.
Although natural for specialists, we initially introduce the result below, which is a consequence of the second resolvent identity [6] and the fact that operators with compact resolvent have purely discrete and unbounded spectrum (see Theorem A.1 in Appendix A).
Proposition 2.1.
Let $A$ be a bounded operator in $\ell^{2}(\mathbb{Z})$ and let $V:\mathbb{Z}\longrightarrow\mathbb{R}$ with
$\lim_{n\to\mp\infty}V(n)=\mp\infty;$
so $H=A+V$ has purely simple discrete spectrum and its eigenvalues satisfy
$\lim_{n\to\mp\infty}\lambda_{n}=\mp\infty.$
Since $\{\lambda_{n}\}$ is discrete, one writes
$\dots,\lambda_{-(n+1)}\leq\lambda_{-n},\dots,\lambda_{-2}\leq\lambda_{-1}\leq\lambda_{0}\leq\lambda_{1}\leq\lambda_{2},\dots,\lambda_{n}\leq\lambda_{n+1},\dots$ (3)
where $\lambda_{0}=\displaystyle\min\{\lambda_{n}\mid\,\lambda_{n}\geq 0\}$.
Remark 2.1.
  1. i)
    Note that with the usual indexing (3), one has that $\{\lambda_{\mp n}\}_{n\in\mathbb{N}}$ are monotonic.
  2. ii)
    Condition $\lim_{n\to\pm\infty}|V(n)|=+\infty$ is important to ensure that $H$ has a purely discrete spectrum. As is well known, e.g., only $\limsup_{n\to\pm\infty}|V(n)|=+\infty$ is not sufficient; see, for example, [16] for cases involving Schrödinger operators with sparse potentials.
  3. iii)
    For the reader’s convenience, we present a proof of this result in Appendix A.
Provided that $V$ grows according to a power-law (or logarithm) order, our next result establishes that the asymptotic behavior of the eigenvalues is determined by the growth of $V$, as long as the eigenfunctions exhibit a semi-uniform polynomial decay.
We shall prove the following
Theorem 2.1.
Let $H=A+V$, where $A$ is bounded and $V:\mathbb{Z}\longrightarrow\mathbb{R}$ with $\lim_{n\to\mp\infty}V(n)=\mp\infty$ such that, for some $\alpha>0$,
$\lim_{n\to\mp\infty}\frac{|V(n)|}{|n|^{\alpha}}=a\not=0\quad or\quad\lim_{n\to\mp\infty}\frac{|V(n)|}{\log(|n|+1)}=a\not=0.$
Suppose that there is an orthonormal basis of $\ell^{2}(\mathbb{Z})$ formed by eigenfunctions of $H$$\{\phi_{n}\}$, that is, $H\phi_{n}=\lambda_{n}\phi_{n}$, for which there exist a $\gamma>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma}{|m-n|},\,m,n\in\mathbb{Z}.$
Then,
$\lim_{n\to\mp\infty}\frac{\lambda_{n}}{V(n)}=1.$
Under additional assumptions, we can obtain the following refinement.
Theorem 2.2.
Let $H=A+V$, where $A$ is bounded and $V$ given by
$V(n)=n+b(n),\,n\in\mathbb{Z},$
for some $b\in\ell^{\infty}(\mathbb{Z})$. Suppose that there is an orthonormal basis of $\ell^{2}(\mathbb{Z})$ formed by eigenfunctions of $H$$\{\phi_{n}\}$, that is, $H\phi_{n}=\lambda_{n}\phi_{n}$, for which there exist a $\gamma>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma}{|m-n|},\,m,n\in\mathbb{Z}.$
Then, there exists a $\gamma^{\prime}>0$ such that,
$|\lambda_{n}-n|\leq\gamma^{\prime},\,n\in\mathbb{Z}.$
Remark 2.2.
  1. i)
    The results stated in Theorems 2.1 and 2.2 are expected in the following sense: consider the eigenvalue equation
    $\lambda_{m}\phi_{m}(n)=(H\phi_{m})(n),m,n\in\mathbb{Z},$
    that is,
    $\lambda_{m}\phi_{m}(n)=(A\phi_{m})(n)+V(n)\phi_{m}(n).$
    Squaring both sides and summing over $|n|\leq|m|$, one has
    $|\lambda_{m}|^{2}\sum_{|n|\leq|m|}|\phi_{m}(n)|^{2}$ $=$ $\sum_{|n|\leq|m|}|(A\phi_{m})(n)|^{2}+2\sum_{|n|\leq|m|}Re\biggr{(}\overline{(A\phi_{m})(n)}V(n)\phi_{m}(n)\biggr{)}$
    $+$ $\sum_{|n|\leq|m|}|V(n)\phi_{m}(n)|^{2}$
    $\leq$ $\sum_{|n|\leq|m|}|(A\phi_{m})(n)|^{2}+2\sum_{|n|\leq|m|}Re\biggr{(}\overline{(A\phi_{m})(n)}V(n)\phi_{m}(n)\biggr{)}$
    $+$ $\max_{|n|\leq|m|}|V(n)|^{2}\sum_{|n|\leq|m|}|\phi_{m}(n)|^{2}.$
    Similarly, one has also
    $|V(n)|^{2}\sum_{|m|\leq|n|}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{|m|\leq|n|}|(A\phi_{m})(n)|^{2}-2\sum_{|m|\leq|n|}Re\biggr{(}\overline{(A\phi_{m})(n)}\lambda_{m}\phi_{m}(n)\biggr{)}$
    $+$ $\max_{|m|\leq|n|}|\lambda_{m}|^{2}\sum_{|m|\leq|n|}|\phi_{m}(n)|^{2}.$
    The inequalities above indicate that, in fact, the asymptotic behavior of the eigenvalues is determined by the growth of $V$, given that $A$ is bounded. However, there is a subtlety: it is necessary to control the sums
    $\sum_{|n|\leq|m|}|\phi_{m}(n)|^{2}\quad\text{and}\quad\sum_{|m|\leq|n|}|\phi_{m}(n)|^{2}$
    to ensure these quantities are not ”small.” For instance, consider $\{\phi_{m}\}_{m\in\mathbb{Z}}$, with $\phi_{m}:=\delta_{m+1}$ (the canonical orthonormal basis shifted by one position). Then, for every $m\in\mathbb{N}$,
    $\sum_{|n|\leq|m|}|\phi_{m}(n)|^{2}=0.$
    In this context, the hypothesis of semi-uniform polynomial decay of the eigenfunctions allows us to resolve this issue (see Propositions 4.1 and 4.2 below) and obtain the result.
  2. ii)
    The main ingredient of the proofs is a relation, established by Propositions 4.1 and 4.2 (see ahead), between the semi-uniform polynomial decay of the eigenfunctions and the manner in which their support ”concentrates” relative to $\ell^{2}(\mathbb{Z}_{+})$ and $\ell^{2}(\mathbb{Z}_{-})$. The arguments presented rely solely on classical functional analysis in Hilbert spaces. In this context, we believe that our work has the potential to appeal to a broad range of readers.
  3. iii)
[…]
  4. iv)
    It is worth underlining that the study of the asymptotic behavior of eigenvalues of operators, for instance, of the Schrödinger-type, in $\ell^{2}(\mathbb{Z}^{d})$ or $L^{2}(\mathbb{R}^{d})$ is a classical subject that plays an important role in various topics in analysis and geometry (see [3, 4] and the references therein). We expect that the techniques presented here may be extended to operators in $L^{2}(\mathbb{R}^{d})$, with the potential to uncover other applications.
2.2 Decay of eigenfunctions
As mentioned in the Introduction, we conclude this section by proving Theorem 1.4.
2.2.1 Proof of Theorem 1.4
Note that $ii)\Rightarrow i)$ follows directly from Theorem 2.2 by taking $A=T_{a}$; thus, it remains to prove $i)\Rightarrow ii)$. Let $\{\phi_{m}\}$ be an orthonormal basis of $\ell^{2}(\mathbb{Z})$ formed by eigenfunctions of $T_{0}+b$, that is, $(T_{0}+b)\phi_{m}=\lambda_{m}\phi_{m},\,m\in\mathbb{Z}$.
Let $r\in\mathbb{N}\cup\{0\}$. By the hypothesis, there exists a $\gamma>0$ such that $|\lambda_{m}-m-b(n)|\leq\gamma$ for all $m,n\in\mathbb{Z}$. Note that for $|m-n|\leq 2\gamma$, one has
$|\phi_{m}(n)|\leq 1\leq\frac{(2\gamma)^{r+1}}{|m-n|^{r+1}}.$
Thus, we need to estimate only the case where $|m-n|>2\gamma$.
Observe that
$((T_{0}+b)\phi_{m})(n)=\lambda_{m}\phi_{m}(n)\iff\phi_{m}(n)=\frac{(T_{a}\phi_{m})(n)}{\lambda_{m}-(n+b(n))},$
which implies that
$\phi_{m}(n)=\frac{\sum_{k\in\mathbb{Z}}a(k)\phi_{m}(n-k)}{(\lambda_{m}-m-b(n))-(n-m)}.$
Thus, we obtain
$|\phi_{m}(n)|\leq\frac{\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{||\lambda_{m}-m-b(n)|-|n-m||}.$
Since
$|\lambda_{m}-m-b(n)|\leq\gamma,\quad m,n\in\mathbb{Z},$
for $|m-n|>2\gamma$, we get
$|\phi_{m}(n)|$$\leq$ $\frac{\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{||\lambda_{m}-m-b(n)|-|n-m||}$
$=$ $\frac{\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{|n-m|-|\lambda_{m}-m-b(n)|}\leq\frac{\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{|n-m|-\gamma},$
Hence,
$|\phi_{m}(n)|\leq\frac{4\gamma\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{|m-n|},\quad|m-n|>2\gamma.$ (4)
Thus, if $r=0$, then there exists $\gamma_{0}^{\prime}>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma_{0}^{\prime}}{|m-n|},\quad|m-n|>2\gamma.$
As previously discussed, this is sufficient to ensure that there exists $\gamma_{0}>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma_{0}}{|m-n|},\quad m,n\in\mathbb{Z}.$ (5)
This proves the case $r=0$. For any $r\in\mathbb{N}$, it suffices to carefully apply inequality (4) recursively. In fact, we now proceed to prove the case $r=1$. Recall that we need to consider only the case $|m-n|>2\gamma$.
First, we apply inequality (4) for $|(m-n)+k|\geq 2$ and $|k|\geq 2$. In this case, we observe that
$|m-n|\leq|(m-n)+k|+|k|\leq|(m-n)+k||k|,$
(recall that if $a,b\geq 2$, then $a+b\leq ab$), so we obtain
$|\phi_{m}(n-k)|\leq\frac{4\gamma\sum_{j\in\mathbb{Z}}|a(j)||\phi_{m}(n-k-j)|}{|(m-n)+k|}\leq\frac{4\gamma|k|}{|m-n|}\sum_{j\in\mathbb{Z}}|a(j)||\phi_{m}(n-k-j)|.$ (6)
Thus, for $|(m-n)+k|\geq 2$ and $|k|\geq 2$, we have
$|\phi_{m}(n-k)|\leq\frac{4\gamma\gamma_{0}|k|}{|m-n|}.$
Since $a(0)=0$, we then obtain
$|\phi_{m}(n)|\leq\frac{4\gamma\sum_{k\in\mathbb{Z}}|a(k)||\phi_{m}(n-k)|}{|m-n|}$$\leq$ $\frac{4\gamma\gamma_{0}}{|m-n|^{2}}\sum_{|k|\geq 2,\,|(m-n)+k|\geq 2}|k||a(k)|$
$+$ $\frac{4\gamma|a(\mp 1)||\phi_{m}(n\pm 1)|}{|m-n|}$
$+$ $\frac{4\gamma|a((n-m))||\phi_{m}(n-(n-m))|}{|m-n|}$
$+$ $\frac{4\gamma|a((n-m)\mp 1)||\phi_{m}(n-(n-m)\pm 1)|}{|m-n|}.$
Finally, for $|m-n|>1$,
$|\phi_{m}(n\pm 1)|\leq\frac{\gamma_{0}}{|m-n\pm 1|}\leq\frac{\gamma_{0}}{||m-n|\mp 1|}\leq\frac{4\gamma_{0}}{|m-n|}.$
Since $a\in\ell^{2}_{1}(\mathbb{Z})$, there exists a $\gamma^{\prime\prime}>0$ such that,
$|a(n-m)|\leq\frac{\gamma^{\prime\prime}}{|m-n|},$
and a $\gamma^{\prime\prime\prime}>0$ such that for $|m-n|>1$,
$|a((n-m)\mp 1)|\leq\frac{\gamma^{\prime\prime\prime}}{|m-n|}.$
Combining these inequalities above with the fact that $|\phi_{m}(n)|\leq 1,\,m,n\in\mathbb{Z}$, we conclude that there exists some $\gamma_{1}$ depending only on $||a||_{1}$ such that
$|\phi_{m}(n)|\leq\frac{\gamma_{1}}{|m-n|^{2}},\quad m,n\in\mathbb{Z}.$ (7)
By applying inequalities (4), (5), (6), and (7) recursively, one obtains the case $r=2$, and so on for higher values of $r$.
3 Power-Law SULE revisited and Proof of Theorem 1.5
Next, we present a proof of Theorem 1.5. The main ingredient of this proof is a notion of Power-Law SULE (Lemma 3.1), which will be combined with Theorem 1.4.
Although it has been known by experts for a long time and discussed in recent papers, mainly by Shi et al. [19, 20, 21] and Sun et al. [15, 26], that a semi-uniform polynomial decay of eigenfunctions of operators in $\ell^{2}(\mathbb{Z})$ with pure point spectrum implies a limitation on moments, the author is not aware of any reference that provides a detailed explanation of a general notion of Power-Law SULE for arbitrary operators in $\ell^{2}(\mathbb{Z})$ (an analysis in this direction applied to polynomial long-range hopping random operators can be found in [15]). In this context, we revisit below a notion of Power-Law SULE for general operators with pure point spectrum acting in $\ell^{2}(\mathbb{Z})$.
Lemma 3.1 (Strong Power-Law SULE).
Let $H$ be a self-adjoint operator acting in $\ell^{2}(\mathbb{Z})$. Suppose that there exists an orthonormal basis of $\ell^{2}(\mathbb{Z})$ by eigenfunctions of $H$, $\{\phi_{m}\}_{m\in\ \mathbb{Z}}$, and an $\alpha>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma_{\alpha}}{|n-m|^{\alpha}},\,m,n\in\mathbb{Z},$
for some $\gamma_{\alpha}>0$ depending only on $\alpha$. If $\alpha>3/2+q/2$, then
$\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-itH}\delta_{k},\delta_{n}\rangle|^{2}<C_{q,k}.$
Remark 3.1.
  1. i)
    Lemma 3.1) show that a semi-uniform polynomial decay of eigenfunctions leads directly to uniform power-law dynamical localization. Although it has been expected and known by experts for a long time that a variation of SULE [11] implies a variation of dynamical localization, at least as far as this author’s knowledge goes, most specialists have recently used an approach based on Sobolev-type norms to obtain power-law dynamical localization through the semi-uniform decay of eigenfunctions [14, 19, 20, 21, 26]. Our approach is more direct and natural, yet little discussed in the literature.
  2. ii)
    As previously mentioned, notion of Strong Power-Law SULE (Lemma 3.1) can also be applied directly to Maryland-type potentials [24] (see Corollary 2.10 in [20]). It is worth noting that in this paper, we revisit a notion of Power-Law SULE that is slightly different from what would be a completely analogous approach to what is treated in [11]; namely, we assume a stronger hypothesis, which serves our objective in this work. It seems possible to obtain a completely analogous version; to do so, one must prove a version of the Theorem 7.1 in [11] in the power-law case and follow the same proof of the Lemma 3.1.
  3. iii)
    The proof of the Lemma 3.1 involves only suitable manipulations, avoiding the singularities of the function $1/|m-\cdot|$ and in fact that for $a,b\geq 2$, one has $a+b\leq ab$.
Proof.
Let $k\in\mathbb{Z}$ and let $\delta_{k}$, $(\delta_{k})(j)=\delta_{jk}$, $k,j\in\mathbb{Z}$. Note that:
$\langle e^{-itH}\delta_{k},\delta_{n}\rangle$ $=$ $\sum_{m\in\mathbb{Z}}\overline{\langle\phi_{m},\delta_{k}\rangle}\langle e^{-itH}\phi_{m},\delta_{n}\rangle$
$=$ $\sum_{m\in\mathbb{Z}}e^{it\lambda_{m}}\overline{\langle\phi_{m},\delta_{k}\rangle}\langle\phi_{m},\delta_{n}\rangle,n\in\mathbb{Z}.$
We consider below $n\neq k,k\mp 1$ and $|n|\geq 2$. Since $|\phi_{m}(n)|\leq 1,\,m,n\in\mathbb{Z}$ and
$|\phi_{m}(n)|\leq\frac{\gamma_{\alpha}}{|n-m|^{\alpha}},\,m,n\in\mathbb{Z},$
one has
$|\langle e^{-itH}\psi,\delta_{n}\rangle|$$\leq$ $\sum_{m\in\mathbb{Z}}|\langle\phi_{m},\delta_{k}\rangle||\langle\phi_{m},\delta_{n}\rangle|=\sum_{m\in\mathbb{Z}}|\phi_{m}(k)||\phi_{m}(n)|$ (8)
$=$ $\biggr{\{}\displaystyle\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|\biggr{\}}$
$+$ $|\phi_{n}(k)||\phi_{n}(n)|+|\phi_{k}(k)||\phi_{k}(n)|$
$+$ $|\phi_{n\mp 1}(k)||\phi_{n\mp 1}(n)|+|\phi_{k\mp 1}(k)||\phi_{k\mp 1}(n)|$
$+$ $|\phi_{-1}(k)||\phi_{-1}(n)|+|\phi_{0}(k)||\phi_{0}(n)|+|\phi_{1}(k)||\phi_{1}(n)|$
$\leq$ $\biggr{\{}\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|\biggr{\}}$
$+$ $\frac{2\gamma_{\alpha}}{|n-k|^{\alpha}}+\frac{\gamma_{\alpha}}{|(n\mp 1)-k|^{\alpha}}+\frac{\gamma_{\alpha}}{|n-(k\mp 1)|^{\alpha}}$
$+$ $\frac{\gamma_{\alpha}}{|n+1|^{\alpha}}+\frac{\gamma_{\alpha}}{|n|^{\alpha}}+\frac{\gamma_{\alpha}}{|n-1|^{\alpha}}.$
One has also
[…]
We analyze two distinct cases:
Case $|k|\geq 2$. for $|n-m|\geq 2$ and $|k-m|\geq 2$, since $a,b\geq 2$ implies $a+b\leq ab$,
$|n-k|$$\leq$ $|k-m|+|n-m|\leq|k-m||n-m|,$
for $|k-m|\geq 2$ and $|k|\geq 2$,
$|m|=|m-k+k|\leq|m-k|+|k|\leq|m-k||k|,$
$\frac{|m|}{|k|}\leq|k-m|\leq|k-m||n-m|.$
Thus, in this case, for every $\epsilon>0,$
$\frac{1}{|k-m|^{\alpha}}\frac{1}{|n-m|^{\alpha}}$ $=$ $\frac{1}{(|n-m||k-m|)^{(\alpha-1-\epsilon/2)+1+\epsilon/2}}$
$=$ $\frac{1}{(|n-m||k-m|)^{\alpha-1-\epsilon/2}}\frac{1}{(|n-m||k-m|)^{1+\epsilon/2}}$
$\leq$ $\frac{1}{|n-k|^{\alpha-1-\epsilon/2}}\frac{|k|^{1+\epsilon/2}}{|m|^{1+\epsilon/2}}$
and, therefore,
$\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|\leq\frac{A_{0}\gamma_{\alpha}^{2}|k|^{1+\epsilon/2}}{|n-k|^{\alpha-1-\epsilon/2}}$
where $A_{0}=\displaystyle\sum_{m\neq 0}\frac{1}{|m|^{1+\epsilon/2}}$. Since $\alpha>\epsilon_{0}+3/2+q/2$ for some $\epsilon_{0}>0$, one has
$\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|$$\leq$ $\frac{A_{0}\gamma_{\alpha}^{2}|k|^{1+\epsilon_{0}/2}}{|n-k|^{\alpha-1-\epsilon_{0}/2}}$ (9)
$\leq$ $\frac{A_{0}\gamma_{\alpha}^{2}|k|^{1+\epsilon_{0}/2}}{|n-k|^{\frac{1+q+\epsilon_{0}}{2}}}.$
Combining (8) with (9), is obtained
$\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-itH}\delta_{k},\delta_{n}\rangle|^{2}<C_{q,k}.$
Case $|k|\leq 1$: for $|n-m|\geq 2$, $|k-m|\geq 2$ and $|m|\geq 2$, since $|n|\geq 2$,
$|n-k|$$\leq$ $|n|+|k|\leq|n|+2\leq 2|n|\leq 2(|m|+|n-m|)$
$\leq$ $2|m||n-m|\leq 2(|m-k|+|k|)|n-m|$
$\leq$ $2(|m-k|+|m-k|)|n-m|=4|m-k||n-m|,$
$|m|-1\leq|m|-|k|\leq|m-k|\leq|m-k||n-m|.$
Thus, in this case, for every $\epsilon>0,$
$\frac{1}{|k-m|^{\alpha}}\frac{1}{|n-m|^{\alpha}}$ $=$ $\frac{1}{(|n-m||k-m|)^{(\alpha-1-\epsilon/2)+1+\epsilon/2}}$
$=$ $\frac{1}{(|n-m||k-m|)^{\alpha-1-\epsilon/2}}\frac{1}{(|n-m||k-m|)^{1+\epsilon/2}}$
$\leq$ $\frac{4^{\alpha-1-\epsilon/2}}{|n-k|^{\alpha-1-\epsilon/2}}\frac{1}{(|m|-1)^{1+\epsilon/2}}$
and, therefore,
$\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|\leq\frac{A_{0}\gamma_{\alpha}^{2}4^{\alpha-1-\epsilon/2}}{|n-k|^{\alpha-1-\epsilon/2}},$
where $A_{0}=\displaystyle\sum_{|m|\geq 2}\frac{1}{(|m|-1)^{1+\epsilon/2}}$. Since $\alpha>\epsilon_{0}+3/2+q/2$ for some $\epsilon_{0}>0$, one has
$\sum_{\begin{subarray}{c}m\not=n\,m\not=n\mp 1,|m|\geq 2\\ m\not=k,\,m\not=k\mp 1\end{subarray}}|\phi_{m}(k)||\phi_{m}(n)|$$\leq$ $\frac{A_{0}\gamma_{\alpha}^{2}4^{\alpha-1-\epsilon_{0}/2}}{|n-k|^{\alpha-1-\epsilon_{0}/2}}$ (10)
$\leq$ $\frac{A_{0}\gamma_{\alpha}^{2}4^{\alpha-1-\epsilon_{0}/2}}{|n-k|^{\frac{1+q+\epsilon_{0}}{2}}}.$
Finally, combining (8) with (10), again is obtained
$\sup_{t\in\mathbb{R}}\sum_{n\in\mathbb{Z}}|n|^{q}|\langle e^{-itH}\delta_{k},\delta_{n}\rangle|^{2}<C_{q}.$
This concludes the proof of the theorem. ∎
To prove Theorem 1.5, we will also use the following results presented in [8, 10, 26].
Theorem 3.1 (Theorem 2.6 in [26]).
To each $r\geq 0$ and $a\in\ell_{r}^{1}(\mathbb{Z})$, $T_{a}$ has purely discrete spectrum $\sigma(T_{a})=\mathbb{Z}$.
Remark 3.2.
Although this result, in this form, was originally presented in [26], the standard argument using Fourier series employed in its proof was initially introduced in [10].
Regarding the persistence of pure point spectrum for perturbations of one-dimensional operators with discrete spectra. Denote by $A_{\tau}$ ($\tau\geq 0$) the Banach space of linear operators $c$ in $\ell^{2}(\mathbb{Z})$ with matrix representation $c=\{c_{nm}\}$ in the orthonormal basis $\{\phi_{m}\}$ of eigenfunctions of $T_{0}$, that is,
$c_{nm}:=\langle\phi_{n},c\phi_{m}\rangle,\quad n,m\in\mathbb{Z},$
and finite norm
$N_{\tau}(c):=\sup_{n}\sum_{m}|c_{nm}|e^{\tau|n-m|}<+\infty.$
The next result is a consequence from the KAM method.
Theorem 3.2 (Theorem 1.1 in [8] or Theorem 4.2 in [10]).
Let $T_{0}+b$, $b\in\ell^{\infty}(\mathbb{Z})$. If $||b||_{\infty}$ is small enough, then there exists a $\tau>0$ and unitary operator $P\in A_{\tau}$ such that
$P^{-1}(T_{0}+b)P$
is diagonal with a discrete and simple spectrum, and whose eigenfunctions are $\{\phi_{m}\}_{m\in\mathbb{Z}}$. Hence, $\{P\phi_{m}\}_{m\in\mathbb{Z}}$ forms an orthonormal basis of $\ell^{2}(\mathbb{Z})$, consisting of eigenfunctions of $T_{0}+b$.
3.1 Proof of Theorem 1.5
We will prove item $ii)$. Since, by the Theorem 3.1, $\sigma(T_{a})=\mathbb{Z}$, then, by the Theorem 1.4, the result follows for $b=0$. Thus, there exists a $\gamma_{r}>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma_{r}}{|n-m|^{r+1}}.$
Note that, for small enough $||b||_{\infty}$, by the Theorem 3.2, it is sufficient to estimate $|(P\phi_{m})(n)|$ with $P\in\mathcal{A}(\tau)$ for some $\tau>0$. There exists a $\gamma^{\prime}>0$ such that
$|P_{km}|\leq\gamma^{\prime}\,\frac{e^{-\frac{\tau|k-m|}{2}}}{|k-m|^{r+1}},\,m\neq k\quad{\rm and}\quad|P_{mm}|\leq\gamma^{\prime}.$
Thus,
$|(P\phi_{m})(n)|=|\langle P\phi_{m},\delta_{n}\rangle|$ $=\left\lvert\sum_{k\in\mathbb{Z}}\overline{\langle\delta_{n},\phi_{k}\rangle}\langle\phi_{k},P\phi_{m}\rangle\right\rvert$
[…]
$+\frac{\gamma_{r}+\gamma^{\prime}}{|n-m|^{r+1}}+\frac{\gamma_{r}}{|n-(m\mp 1)|^{r+1}}+\frac{\gamma^{\prime}}{|m-(n\mp 1)|^{r+1}}$
[…]
$+\frac{\gamma_{r}+\gamma^{\prime}}{|n-m|^{r+1}}+\frac{\gamma_{r}}{|n-(m\mp 1)|^{r+1}}+\frac{\gamma^{\prime}}{|m-(n\mp 1)|^{r+1}}.$
Since for $|m-k|\geq 2$ and $|n-k|\geq 2$,
$|n-m|\leq|m-k|+|n-k|\leq|m-k||n-k|,$
one has
$|(P\phi_{m})(n)|$$\leq$ $\sum_{k\in\mathbb{Z}}e^{-\frac{\tau|k|}{2}}\frac{\gamma_{r}}{|n-m|^{r+1}}$
$+$ $\frac{\gamma_{r}+\gamma^{\prime}}{|n-m|^{r+1}}+\frac{\gamma_{r}}{|n-(m\mp 1)|^{r+1}}+\frac{\gamma^{\prime}}{|m-(n\mp 1)|^{r+1}}.$
This concludes the proof of item $ii)$ and also of the Theorem 1.5. In fact, $i)$ follows from $ii)$ and Theorem 1.4, while $iii)$ follows from the Lemma 3.1. ∎
4 Preparation for the proofs of Theorems 2.1 and 2.2
4.1 Properties of semi-uniform polynomial decay
Our next technical result, which will be used in the proof of Theorems 2.1 and 2.2, establishes a concentration property for an orthonormal basis of $\ell^{2}(\mathbb{Z})$, under the assumption of semi-uniform polynomial decay.
Proposition 4.1.
Let $\{\phi_{m}\}$ be an orthonormal basis of $\ell^{2}(\mathbb{Z})$. If there exists $\gamma>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma}{|m-n|},$
then, for each $10<\delta<$ 1, there exists a fixed $N>0$, that does not depend on $m$, such that
$\sum_{|n|\leq|m|+N}|\phi_{m}(n)|^{2}\geq 1-\delta,\quad\forall m\in\mathbb{Z}.$
Proof.
Fix $10<\delta<$ 1. Since
$\sum_{n\in\mathbb{Z}}|\phi_{m}(n)|^{2}=1,\quad\forall m\in\mathbb{Z}.$
We need to show that there exists $N>0$, independent of $m$, such that
$\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}\leq\delta.$
As $|\phi_{m}(n)|\leq\displaystyle\frac{\gamma}{|m-n|}$, one has
$\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}\leq\sum_{|n|>|m|+N}\frac{\gamma^{2}}{|m-n|^{2}}.$
Sinse $||n|-|m||\leq|m-n|$. Making $k=|n|-|m|$, one has also
$\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}$$\leq$ $\gamma^{2}\sum_{|n|>|m|+N}\frac{1}{|m-n|^{2}}=\gamma^{2}\sum_{|n|-|m|>N}\frac{1}{|m-n|^{2}}$
$\leq$ $\gamma^{2}\sum_{|n|-|m|>N}\frac{1}{||n|-|m||^{2}}=\gamma^{2}\sum_{k>N}\frac{1}{k^{2}}$
[…]
Thus,
$\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}\leq\frac{\gamma^{2}}{N}.$
Therefore, it is sufficient to choose $N>\lceil\gamma^{2}/\delta\rceil$, and it follows that
$\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}\leq\delta.$
Hence,
$\sum_{|n|\leq|m|+N}|\phi_{m}(n)|^{2}=1-\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}\geq 1-\delta,\quad\forall m\in\mathbb{Z}.$
Our next result, which will also be important for proving Theorems 2.1 and 2.2, can be interpreted as follows: under the assumption of semi-uniform polynomial decay, the eigenfunctions $\phi_{m}$ corresponding to positive indices are asymptotically concentrated in $\ell^{2}(\mathbb{Z}_{+})$, while those corresponding to negative indices are asymptotically concentrated in $\ell^{2}(\mathbb{Z}_{-})$.
Proposition 4.2.
Let $\{\phi_{m}\}$ be an orthonormal basis of $\ell^{2}(\mathbb{Z})$. If there exists $\gamma>0$ such that
$|\phi_{m}(n)|\leq\frac{\gamma}{|m-n|},$
then, for every each $10<\delta<$ 1, there exists a fixed $M>0$, that does not depend on $n$, such that
$\sum_{m=0}^{\mp(n+M)}|\phi_{m}(\mp n)|^{2}\geq 1-\delta-\frac{\gamma^{2}}{n},\quad\forall n\in\mathbb{N}.$
Proof.
Fix $10<\delta<$ 1 and let $n\in\mathbb{N}$. Arguing identically to how it was done in the proof of the Proposition 4.1, it follows that for $M>\lceil\gamma^{2}/\delta\rceil$
$\sum_{m=0}^{-(n+M)}|\phi_{m}(n)|^{2}+\sum_{|m|>n+M}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{m=0}^{-(n+M)}|\phi_{m}(n)|^{2}+\delta.$
Since $n>0$ and $m\leq 0$, $|n-m|=n+|m|$. Thus,
$\sum_{m=0}^{-(n+M)}|\phi_{m}(n)|^{2}+\sum_{|m|>n+M}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{m=0}^{-(n+M)}|\phi_{m}(n)|^{2}+\delta\leq\gamma^{2}\sum_{m=0}^{-(n+M)}\frac{1}{|n-m|^{2}}+\delta$
$=$ $\gamma^{2}\sum_{m=0}^{-(n+M)}\frac{1}{(n+|m|)^{2}}+\delta$
$=$ $\gamma^{2}\sum_{m=0}^{n+M}\frac{1}{(n+m)^{2}}+\delta=\gamma^{2}\sum_{k=n}^{2n+M}\frac{1}{k^{2}}+\delta$
$\leq$ $\gamma^{2}\sum_{k=n}^{\infty}\frac{1}{k^{2}}+\delta\leq\frac{\gamma^{2}}{n}+\delta.$
Hence,
$\sum_{m=0}^{(n+M)}|\phi_{m}(n)|^{2}$ $=$ $1-\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}-\sum_{m=0}^{-(n+M)}|\phi_{m}(n)|^{2}$
$\geq$ $1-\delta-\frac{\gamma^{2}}{n},\quad n\in\mathbb{N}.$
Again, note that for $M>\lceil\gamma^{2}/\delta\rceil$
$\sum_{m=0}^{n+M}|\phi_{m}(-n)|^{2}+\sum_{|m|>n+M}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{m=0}^{n+M}|\phi_{m}(-n)|^{2}+\delta.$
Since $m\geq 0$, one has also
$\sum_{m=0}^{n+M}|\phi_{m}(-n)|^{2}+\sum_{|m|>n+M}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{m=0}^{n+M}|\phi_{m}(-n)|^{2}+\delta\leq\gamma^{2}\sum_{m=0}^{n+M}\frac{1}{|m-(-n)|^{2}}+\delta$
$=$ $\gamma^{2}\sum_{m=0}^{n+M}\frac{1}{(m+n)^{2}}+\delta=\gamma^{2}\sum_{k=n}^{2n+M}\frac{1}{k^{2}}+\delta$
$\leq$ $\gamma^{2}\sum_{k=n}^{\infty}\frac{1}{k^{2}}+\delta\leq\frac{\gamma^{2}}{n}+\delta.$
Hence,
$\sum_{m=0}^{-(n+M)}|\phi_{m}(-n)|^{2}$ $=$ $1-\sum_{|n|>|m|+N}|\phi_{m}(n)|^{2}-\sum_{m=0}^{n+M}|\phi_{m}(-n)|^{2}$
$\geq$ $1-\delta-\frac{\gamma^{2}}{n},\quad n\in\mathbb{N}.$
This concludes the proof of the proposition. ∎
5 Proofs of Theorems 2.1 and 2.2
Next, we present proofs of Theorems 2.1 and 2.2. As already mentioned, the main ingredient of these proofs is a relation, given by Propositions 4.1 and 4.2, between the semi-uniform polynomial decay of the eigenfunctions and the manner in which their support ”concentrates” relative to $\ell^{2}(\mathbb{Z}_{+})$ and $\ell^{2}(\mathbb{Z}_{-})$.
Lemma 5.1.
Let $f:\mathbb{Z}\longrightarrow\mathbb{R}$ be a function such that, for some $\alpha>0$,
$\lim_{n\to\mp\infty}\frac{|f(n)|}{|n|^{\alpha}}=a\not=0\quad or\quad\lim_{n\to\mp\infty}\frac{|f(n)|}{\log(|n|+1)}=a\not=0.$
Then, for every $N\in\mathbb{N}$,
$\lim_{n\to\mp\infty}\frac{\displaystyle\max_{|m|\leq|n|+N}|f(m)|}{|f(n)|}=1.$
Proof.
First case:
$\lim_{n\to\mp\infty}\frac{|f(n)|}{|n|^{\alpha}}=a\not=0;$
so, for every $\epsilon>0$ and for all $n>n_{\epsilon}$,
$\frac{\displaystyle\max_{|m|\leq|n|+N}|f(m)|}{|f(n)|}$ $=$ $\frac{\displaystyle\max_{|m|\leq|n|+N}|f(m)|}{||n|+N|^{\alpha}}\frac{||n|+N|^{\alpha}}{|n|^{\alpha}}\frac{|n|^{\alpha}}{|f(n)|}$
$\leq$ $a(1+\epsilon)\frac{||n|+N|^{\alpha}}{|n|^{\alpha}}\frac{|n|^{\alpha}}{|f(n)|},$
$\frac{\displaystyle\max_{|m|\leq|n|+N}|f(m)|}{|f(n)|}$$\geq$ $a(1-\epsilon)\frac{||n|+N|^{\alpha}}{|n|^{\alpha}}\frac{|n|^{\alpha}}{|f(n)|}.$
Thus, making $n\to\mp\infty$, one has
$(1-\epsilon)\leq\lim_{n\to\mp\infty}\frac{\displaystyle\max_{|m|\leq|n|+N}|f(m)|}{|f(n)|}\leq(1+\epsilon).$
As $\epsilon>0$ is arbitrary, the result follows. The proof for the second case relies on the same arguments presented above and is therefore omitted. ∎
5.1 Proof of Theorem 2.1
We begin with the following eigenvalue equation, $m,n\in\mathbb{Z},$
$\lambda_{m}\phi_{m}(n)=(H\phi_{m})(n),$
that is,
$\lambda_{m}\phi_{m}(n)=(A\phi_{m})(n)+V(n)\phi_{m}(n).$
Squaring both sides, one has
$|\lambda_{m}\phi_{m}(n)|^{2}$ $=$ $|(A\phi_{m})(n)|^{2}+2Re\biggr{(}\overline{(A\phi_{m})(n)}V(n)\phi_{m}(n)\biggr{)}$
$+$ $|V(n)\phi_{m}(n)|^{2}.$
Thus,
$|\lambda_{m}|^{2}|\phi_{m}(n)|^{2}$$\leq$ $|(A\phi_{m})(n)|^{2}+2|V(n)||(A\phi_{m})(n)||\phi_{m}(n)|+|V(n)||\phi_{m}(n)|^{2}.$
Let $10<\delta<$ 1 and let $N$ (depending on $\delta$) given by Proposition 4.1, summing over $|n|\leq|m|+N$,
$|\lambda_{m}|^{2}\sum_{|n|\leq|m|+N}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)|^{2}$ (11)
$+$ $2\max_{|n|\leq|m|+N}|V(n)|$
$\times$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)||\phi_{m}(n)|$
$+$ $\max_{|n|\leq|m|+N}|V(n)|^{2}\sum_{|n|\leq|m|+N}|\phi_{m}(n)|^{2}.$
As $||\phi_{m}||_{2}=1$, by the choice of $N$ and by the Cauchy-Schwartz inequality, so
$(1-\delta)|\lambda_{m}|^{2}$$\leq$ $||A||^{2}+2||A||\max_{|n|\leq|m|+N}|V(n)|+\max_{|n|\leq|m|+N}|V(n)|^{2}.$
Therefore, for all $m$ large enough,
$\frac{(1-\delta)|\lambda_{m}|^{2}}{|V(m)|^{2}}$$\leq$ $\frac{||A||^{2}}{|V(m)|^{2}}+\frac{2||A||\displaystyle\max_{|n|\leq|m|+N}|V(n)|}{|V(m)|^{2}}+\frac{\displaystyle\max_{|n|\leq|m|+N}|V(n)|^{2}}{|V(m)|^{2}}.$
Finally, since, for some $\alpha>0$,
$\lim_{m\to\mp\infty}\frac{|V(m)|}{|m|^{\alpha}}=a\not=0\quad or\quad\lim_{m\to\mp\infty}\frac{|V(m)|}{\log(|m|+1)}=a\not=0,$
so, by the Lemma 5.1,
$\lim_{m\to\mp\infty}\frac{||A||^{2}}{|V(m)|^{2}}+\frac{2||A||\displaystyle\max_{|n|\leq|m|+N}|V(n)|}{|V(m)|^{2}}+\frac{\displaystyle\max_{|n|\leq|m|+N}|V(n)|^{2}}{|V(m)|^{2}}=1.$
Hence,
$\limsup_{m\to\mp\infty}\frac{(1-\delta)|\lambda_{m}|^{2}}{|V(m)|^{2}}\leq 1.$
Sinse $10<\delta<$ 1 is arbitrary, follows that
$\limsup_{m\to\mp\infty}\frac{|\lambda_{m}|^{2}}{|V(m)|^{2}}\leq 1.$ (12)
Now, let be the subsequence $\lambda_{n}\to+\infty$ as $n\to+\infty$. Again, we consider the eigenvalue equation, $n\in\mathbb{N}$, $m\in\mathbb{N}$
$-V(n)\phi_{m}(n)=(A\phi_{m})(n)-\lambda_{m}\phi_{m}(n).$
Squaring both sides, one has
$|V(n)\phi_{m}(n)|^{2}$ $=$ $|(A\phi_{m})(n)|^{2}$
-- $2Re\biggr{(}\overline{(A\phi_{m})(n)}\lambda_{m}\phi_{m}(n)\biggr{)}$
$+$ $|\lambda_{m}\phi_{m}(n)|^{2}.$
Thus,
$|V(n)\phi_{m}(n)|^{2}$$\leq$ $|(A\phi_{m})(n)|^{2}$
$+$ $2|\lambda_{m}\phi_{m}(n)||(A\phi_{m})(n)|$
$+$ $|\lambda_{m}\phi_{m}(n)|^{2}.$
Let $10<\delta<$ 1 and let $M$ (depending only on $\delta$) given by the Proposition 4.2, summing over $0\leq m\leq n+M$, since
$||\phi_{m}||_{2}=\sum_{m\in\mathbb{Z}}|\phi_{m}(n)|^{2}=\sum_{m\in\mathbb{Z}}|\langle\phi_{m},\delta_{n}\rangle|^{2}=||\delta_{n}||_{2}=1,$
by the same arguments presented above, for every $n$,
$\biggr{(}1-\delta-\frac{\gamma^{2}}{n}\biggr{)}|V(n)|^{2}$$\leq$ $||A||^{2}+2||A||\max_{0\leq m\leq n+M}\lambda_{m}+\max_{0\leq m\leq n+M}\lambda_{m}^{2}.$
Let $\epsilon>0$, by the Lemma 5.1 and (12), for all $n>n_{\epsilon,M}$,
$\biggr{(}1-\delta-\frac{\gamma^{2}}{n}\biggr{)}(1-\epsilon)|V(n+M)|^{2}$$\leq$ $(1-\delta)\frac{|V(n)|^{2}}{|V(n+M)|^{2}}|V(n+M)|^{2}$
$\leq$ $||A||^{2}+2||A||\max_{0\leq m\leq n+M}|V(m)|(1+\epsilon)+\max_{0\leq m\leq n+M}\lambda_{m}^{2}.$
Since $\{\lambda_{n}\}_{n\in\mathbb{N}}$ is monotonic (recall Remark 2.2), making $k=n+M$, for all $k>k_{\epsilon,M}=n_{\epsilon,M}+M$,
$\biggr{(}1-\delta-\frac{\gamma^{2}}{(k-M)}\biggr{)}(1-\epsilon)|V(k)|^{2}$$\leq$ $2||A||^{2}+||A||\max_{0\leq m\leq k}|V(m)|(1+\epsilon)+\max_{0\leq m\leq k}\lambda_{m}^{2}$
$=$ $2||A||^{2}+||A||\max_{0\leq m\leq k}|V(m)|(1+\epsilon)+\lambda_{k}^{2}.$
Thus,
$\liminf_{k\to+\infty}\frac{\lambda_{k}^{2}}{|V(k)|^{2}}\geq(1-\delta)(1-\epsilon).$
Since $10<\epsilon,\delta<$ 1 are arbitrary,
$\liminf_{k\to+\infty}\frac{\lambda_{k}^{2}}{|V(k)|^{2}}\geq 1.$
Finally, note also that the same argument above can be applied identically to the subsequence $\lambda_{n}\to-\infty$ as $n\to-\infty$; so,
$\liminf_{n\to\mp\infty}\frac{|\lambda_{n}|^{2}}{|V(n)|^{2}}\geq 1.$
This concludes the proof of theorem. ∎
5.2 Proof of Theorem 2.2
Let us verify that
$\sup_{n\in\mathbb{Z}_{+}}|\lambda_{n}-n|<+\infty.$
The proof that
$\sup_{n\in\mathbb{Z}_{-}}|\lambda_{n}-n|<+\infty$
is analogous and is omitted for brevity. Let $\delta=\displaystyle\frac{1}{2}$ and let $N$ given by Proposition 4.1, by (11) in the proof of Theorem 2.1, for every $m\in\mathbb{N}$,
$|\lambda_{m}|^{2}\sum_{|n|\leq m+N}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)|^{2}$
$+$ $2\max_{|n|\leq m+N}|V(n)|$
$\times$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)||\phi_{m}(n)|$
$+$ $\max_{|n|\leq m+N}|V(n)|^{2}\sum_{|n|\leq m+N}|\phi_{m}(n)|^{2}.$
As $c=\displaystyle\sum_{|n|\leq m+N}|\phi_{m}(n)|^{2}\geq\frac{1}{2}$, dividing both sides of the equation above by $c$, one has
$\lambda_{m}^{2}$$\leq$ $2\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)|^{2}$
$+$ $4\max_{|n|\leq m+N}|V(n)|$
$\times$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)||\phi_{m}(n)|$
$+$ $\max_{|n|\leq m+N}V(n)^{2}.$
Thus,
$\lambda_{m}^{2}$$\leq$ $2||A||^{2}+4||A||(m+N)+4||A||||b||_{\infty}$
$+$ $\max_{|n|\leq m+N}(n^{2}+2nb(n)+b(n)^{2})$
$\leq$ $2||A||^{2}+4||A||(m+N)+4||A||||b||_{\infty}$
$+$ $(m+N)^{2}+2(m+N)||b||_{\infty}+||b||_{\infty}^{2}.$
Therefore,
$\lambda_{m}^{2}-m^{2}$$\leq$ $2||A||^{2}+4||A||(m+N)+4||A||||b||_{\infty}$
$+$ $2mN+N^{2}+2(m+N)||b||_{\infty}+||b||_{\infty}^{2},$
that is,
$(\lambda_{m}-m)(\lambda_{m}+m)$$\leq$ $2||A||^{2}+4||A||(m+N)+4||A||||b||_{\infty}$
$+$ $2mN+N^{2}+2(m+N)||b||_{\infty}+||b||_{\infty}^{2}.$
Since $\lim_{m\to\infty}\lambda_{m}=+\infty,$ let $m_{0}$ be such that for $m>m_{0}$, $\lambda_{m}>0$; for every $m>m_{0}$,
$\lambda_{m}-m$$\leq$ $\frac{2||A||^{2}+4||A||(m+N)+4||A||||b||_{\infty}}{\lambda_{m}+m}$
$+$ $\frac{2mN+N^{2}+2(m+N)||b||_{\infty}+||b||_{\infty}^{2}}{\lambda_{m}+m}<C.$
for some $C>0$. Thus,
$\lambda_{m}-m\leq\max_{0\leq m\leq m_{0}}(\lambda_{m}-m)+C=:C^{\prime},m\in{\mathbb{Z}}_{+}.$
Now, let $\delta=\frac{1}{2}$ and let $M$ given by Proposition 4.2; argument in a manner analogous to that used to obtain (11), for each $n\in\mathbb{N}$, one has:
$|V(n)|^{2}\sum_{0\leq m\leq n+M}|\phi_{m}(n)|^{2}$$\leq$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)|^{2}$
$+$ $2\max_{0\leq m\leq n+M}|\lambda_{m}|$
$\times$ $\sum_{n\in\mathbb{Z}}|(A\phi_{m})(n)||\phi_{m}(n)|$
$+$ $\max_{0\leq m\leq n+M}|\lambda_{m}|^{2}\sum_{0\leq m\leq n+M}|\phi_{m}(n)|^{2}.$
As $d=\displaystyle\sum_{0\leq m\leq n+M}|\phi_{m}(n)|^{2}\geq\displaystyle\frac{1}{2}-\displaystyle\frac{\gamma^{2}}{n}$, dividing both sides of the equation above by $d$, since $\{\lambda_{n}\}$ is monotonic (recall in Remark 2.1), one has
$|V(n)|^{2}$$\leq$ $2||A||^{2}\frac{1}{\frac{1}{2}-\frac{\gamma^{2}}{n}}+2||A||\,|\lambda_{n+N}|\frac{1}{\frac{1}{2}-\frac{\gamma^{2}}{n}}+\lambda_{n+N}^{2}.$
Let $n_{0}$ be sufficiently large such that for all $n>n_{0}$,
$\frac{1}{2}-\frac{\gamma^{2}}{n}\geq\frac{1}{4}$
and $\lambda_{n}>0$; so, for $n>n_{0}$,
$V(n)^{2}$$\leq$ $8||A||^{2}+8||A||\lambda_{n+M}+\lambda_{n+M}^{2}.$
Making $j=n+M$, one has that for all $j>j_{0}=n_{0}+M$
$V(j-M)^{2}$$\leq$ $8||A||^{2}+8||A||\lambda_{j}+\lambda_{j}^{2},$
that is,
$j^{2}-2jM+M^{2}+2(j-M)b(j-M)+b(j-M)^{2}$$\leq$ $8||A||^{2}+8||A||\lambda_{j}+\lambda_{j}^{2}.$
Thus, for all $j>j_{0}$
$j^{2}-\lambda_{j}^{2}$$\leq$ $8||A||^{2}+8||A||\lambda_{j}$
$+$ $2jM+2(j-M)||b||_{\infty}+||b||_{\infty}^{2},$
which implies that
$j-\lambda_{j}$$\leq$ $\frac{8||A||^{2}+8||A||\lambda_{j}}{j+\lambda_{j}}+\frac{2jM+2(j-M)||b||_{\infty}+||b||_{\infty}^{2}}{j+\lambda_{j}}<C^{\prime\prime}$
for algum $C^{\prime\prime}>0$. Hence,
$j-\lambda_{j}\leq\max_{0\leq j\leq j_{0}}(j-\lambda_{j})+C^{\prime\prime}=:C^{\prime\prime\prime},j\in{\mathbb{Z}}_{+}.$
Finally, let $\gamma^{\prime}:=\max\{C^{\prime},C^{\prime\prime\prime}\}$, then
$\sup_{n\in\mathbb{Z}_{+}}|\lambda_{n}-n|<\gamma^{\prime}.$
Appendix A Discrete spectrum
As mentioned in the introduction, in this appendix, we prove that self-adjoint operators of the form $H=A+V$ in $\ell^{2}(\mathbb{Z})$, where $A$ is bounded and $V$ is such that $\lim_{m\to\mp\infty}V(m)=\mp\infty$, has purely simple discrete spectrum and its eigenvalues satisfy $\lim_{m\to\mp\infty}\lambda_{m}=\mp\infty$ (Proposition 2.1).
We will use the following result.
Theorem A.1 (Theorem 11.3.13. in [6]).
Let $H$ be a self-adjoint operator on $\ell^{2}(\mathbb{Z})$. The following statements are equivalent:
  1. (a)
    There exists an orthonormal basis $(\phi_{m})$ of $\ell^{2}(\mathbb{Z})$ consisting of eigenfunctions of $H$, that is, $H\phi_{m}=\lambda_{m}\phi_{m}$, where the eigenvalues $\lambda_{m}$ are real and form a discrete set. Furthermore, each one of these eigenvalues has finite multiplicity, and $\lim_{m\to\mp\infty}|\lambda_{m}|=+\infty$.
  2. (b)
    $\left(H-zI\right)^{-1}$ is a compact operator for some $z\in\rho(T)$ (and thus for all $z\in\rho(T))$.
A.1 Proof of Proposition 2.1
The result is consequence of the second resolvent identity. Initially, we closely followed the proof of Theorem 4.1 in [10]. Namely, let $\{\delta_{m}\}$ the canonical basis of $\ell^{2}(\mathbb{Z})$ and note that, for every $m\in\mathbb{Z}$ and every $n\in\mathbb{Z}$,
$[V(\delta_{m})](n)=V(n)(\delta_{m})(n)=V(n)(\delta_{nm})=V(m)\delta_{nm}=(V(m)\delta_{m})(n),$
so $V$ satisfies the hypotheses of the Theorem A.1 and, therefore, $\left(V-iI\right)^{-1}$ is a compact operator.
Let us apply the second resolvent identity with:
  • $T=H=A+V$,
  • $S=V$,
  • $z=i$.
Recall the second resolvent identity states:
$\left(T-iI\right)^{-1}-\left(S-iI\right)^{-1}=\left(T-iI\right)^{-1}(S-T)\left(S-iI\right)^{-1}.$
Substituting $T=H=A+V$, $B=V$ and $z=i$, one has
$\left(H-iI\right)^{-1}-\left(V-iI\right)^{-1}=\left(H-iI\right)^{-1}\left(-A\right)\left(V-iI\right)^{-1}.$
Thus,
$\left(H-iI\right)^{-1}=\left(V-iI\right)^{-1}+\left(H-iI\right)^{-1}\left(-V\right)\left(V-iI\right)^{-1}.$
Since $H$ is self-ajoint, $\left(H-iI\right)^{-1}$ is bounded since that in this case $\sigma(H)\subset\mathbb{R}$ [6]. Thus, the sum above is a sum of compact operators and is, therefore, compact. Hence, by Theorem A.1, $H$ has purely simple discrete spectrum and its eigenvalues satisfy
$\lim_{m\to\mp\infty}|\lambda_{m}|=+\infty.$
Finally, let $(\phi_{m})$ be an orthonormal basis consisting of eigenfunctions of $H$. Note that: given $a>0$, if $\lambda_{m}\geq-a$ or $\lambda_{m}\leq a$ for all $m\in\mathbb{N}$. Since, for every $n\in\mathbb{Z}$,
$(A\delta_{n})(n)+V(n)=\langle H\delta_{n},\delta_{n}\rangle=\sum_{m\in\ \mathbb{Z}}\lambda_{n}|\langle\delta_{n},\phi_{m}\rangle|^{2},$
then
$(A\delta_{n})(n)+V(n)\geq-a||\delta_{n}||^{2}=-a\,\,{\rm or}\,\,(A\delta_{n})(n)+V(n)\leq a||\delta_{n}||^{2}=a.$
but this contradicts the fact that $\lim_{n\to\mp\infty}V(n)=\mp\infty$ since that $A$ is bounded. Hence, one has
$\lim_{m\to\mp\infty}\lambda_{m}=\mp\infty.$
Acknowledgments
M. Aloisio thanks the partial support by FAPEMIG (Minas Gerais state agency; under contract 01/24/APQ-03132-24).